RC Meeting and Run Plan, Sunday, March 1 2020

From HallCWiki
Jump to: navigation, search

RC Meeting Information

RC Daily Meetings are M-F at 4:15 PM, 2nd floor Counting House Meeting Room Bluejeans Connection

Sunday 1 Mar 2020.

General updates

  • Beam has been (relatively) quite good the last 24 hours, trip rate has improved
  • Target has been running smoothly in longitudinal at around 53%

Notes for the Future

  • Beam studies to understand what is going on with the machine are planned for Tuesday DAY shift
    • We will try to get in Tuesday morning (RC will talk to PD and RadCon -- may need to open target enclosure)
    • Details being worked out, but it'll be a full 7--8 hour beam studies shift
    • May need to get into the Target enclosure (RadCon support)
    • There are portions of the Beam Studies program that would involve beam delivery into Hall C (no target), so limited opportunity for Access only
    • EPR and pNMR measurements/work are planned during this downtime
  • Injector spot move has been pushed off for at least another week (2--3 hour down)
    • Hope is to push it off to the Hall C down that starts March 13, and then just do a reactivation that would carry us through May 6
      • Not sure how it will work out, but goal will be to minimized downs and maximize integrated current to Hall(s)
  • 'Filler' runs for next spin up / target-down period:
    • 1 hour run with target out for measure of the background just from the beryllium windows

Pending Issues

  • Transverse NMR measurement that failed [1]
  • pNMR issues
  • EPR problems [2]
    • Going forward for EPR, I'd try (1) checking the continuity of the cable leading from the RF amplifier to the EPR coil to see if it's stable, (2) measure the current going through the EPR coil, (3) cleaning the single D2 filter (it was very dirty) and confirming the APD gain is indeed set to maximum, measure the DC light output (currently a measly 3.6 mV), then trying another FM sweep. If that doesn't improve it, then (4) remove the D2 filter, measure the DC light output, perform another FM scan.
  • Kepco power supply commissioning/installation (Todd, Bill)
  • Update Transverse NMR calibration coeff.
  • Persistent BLM & ion chamber trips
    • General activation, weird BLM issues (https://logbooks.jlab.org/entry/3792326), Ad-hoc ion chamber threshold adjustments (https://logbooks.jlab.org/entry/3792095)
    • NOTE: Shift crew should NOT give authorization to MCC change ion chamber thresholds. If they ask, tell Operator to contact an appropriate expert. Make a log entry if trips persist, or call the RC if it becomes severe (for example, >20 trips/hour is getting intolerable)

Opportunistic Access Jobs

  • Target work during Beam Studies (see above)
    • pNMR troubleshooting
    • EPR issues [3]
    • Transverse NMR?[4]

Target coil rotation planned for March 13--20

    • Nominal 1-week downtime in Hall C to rotate target coils, install new cell, and transition to d2n program
    • Status Update for target cell Austin:
      • Wall and window thicknesses done
      • Only 'Reduced set' of density measurements available now, will continue measurements this week (issues with laser at UVa?)
      • It would be best for it to arrive at JLab on March 10 (at the latest) to allow locals to mount cell in oven plate before they get involved with target work in Hall.
    • Lots of preplanning to be completed (RadCon, Techs, S&A, Target group, etc)
    • RW-II training for JP, Arun, Brad (Complete the training documentation and take the RW-II test if you haven't already done so)

Current Run Plan

Sunday DAY → Monday OWL

  • Take production data (HMS/SHMS -3.5/-3.4 GeV) with 3He target, longitudinal polarization with NMR every 5 hrs.

Monday DAY → Tuesday OWL

  • Complete the last ~5 hour NMR run period from OWL, then do this (should happen before noon!)
    • Take longitudinal NMR
    • Rotate target to 90deg Transverse (Target Expert only)
    • Take transverse NMR (If operational, Target expert should know or contact JP)
    • Take 5 hours of production data (HMS/SHMS -3.5/-3.4 GeV) with 3He target in transverse
    • Take transverse NMR (If operational, Target expert should know or contact JP)
    • Rotate target to 180deg Longitudinal (Target Expert only)
    • Take longitudinal NMR
  • Continue production data (HMS/SHMS -3.5/-3.4 GeV) with 3He target, longitudinal polarization with NMR every 5 hrs.

Tuesday DAY

  • Down for beam studies for most of DAY shift
    • Hall C will be in Controlled Access (likely in Laser CANS mode) for part of it
    • Hall C will be locked up for beam studies plans for part of this shift
      • NOTE: Move target to 'NO TARGET' position and be aware of MCC operations when in Beam Permit!

Tuesday SWING → Wed DAY

  • Follow the Short beam recovery procedure (skip the YAG unless you are concerned about the harp scan data)
  • Take production data (HMS/SHMS -3.5/-3.4 GeV) with 3He target, longitudinal polarization with NMR every 5 hrs.

Wednesday DAY → Thursday OWL

  • Complete the last ~5 hour NMR run period from OWL, then do this (should happen before noon!)
    • Take longitudinal NMR
    • Rotate target to 90deg Transverse (Target Expert only)
    • Take transverse NMR (Only if operational, Target expert should know or contact JP)
    • Take 5 hours of production data (HMS/SHMS -3.5/-3.4 GeV) with 3He target in transverse
    • Take transverse NMR (Only if operational, Target expert should know or contact JP)
    • Rotate target to 180deg Longitudinal (Target Expert only)
    • Take longitudinal NMR
  • Continue production data (HMS/SHMS -3.5/-3.4 GeV) with 3He target, longitudinal polarization with NMR every 5 hrs.

Standing Orders

 If beam goes away for more than 2h, or if during the downtime, MCC does operations
 that may change the beam tuning, then do [https://hallcweb.jlab.or/wiki/index.php/Beam_and_Raster_Checkout_for_A1n/d2n_(Short_Procedure)
 beam recovery procedure] (do YAG viewer part only if you are suspicious of
 the harp scan results)
  • NOTE: Shift crew should NOT give authorization to MCC change ion chamber thresholds. If they ask, tell the Operator to contact an appropriate expert. Make a log entry if trips persist, or call the RC if it becomes severe (for example, >20 trips/hour is getting intolerable)
  • Check the position locks are on.
  • Do your best to accommodate Hall A's request for a PITA scan promptly. It is OK to stop the present runs and start new ones in 5--10 minutes after they are done.
  • Check that CODA is running normally (had one strange hangup lately)
  • Check online replays shortly after starting runs, & full replay results too.
  • Make sure all the beam specs are followed (e.g. no raster size below 4mm on 3He cell).
  • Always take 120Hz of EDTM
  • Fill BTA with realistic numbers. You may follow the guidance coming from "load from EPICS", if it does not overestimate the ABU (as it generally does). https://logbooks.jlab.org/entry/3783171. If we have to lower the beam current we actually want because the beam is too unstable, prorate the ABUs.
!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!
!!                                                     !!
!!  ATTN: SHIFT LEADERS                                !!
!!     Please follow the shift summary template on     !!
!!     the laminated SL instructions.                  !!
!!                                                     !!
!!  ATTN: TARGET OPERATORS                             !!
!!  All NMR measurement HCLOG entries should include:  !!
!!     1) A title that clearly lists polarization      !!
!!     2) Screenshots of both NMR_FieldSweep and       !!
!!        the polarization fit                         !!
!!                                                     !!
!!     For an example see the following entry:         !!
!!     https://logbooks.jlab.org/entry/3768696       !!
!!                                                     !!
!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

Notices in Effect for Shift Crew

  • Policy for CANS-related hall status change: Make a logbook entry when the hall goes in, or comes out of laser-controlled access.
  • Magnet policies
    • Spectrometer GUI (go_magnets) should always be visible on Shift Leader computer, hcdesk3
    • HMS Q1 must be ramped down in 50A steps in all cases
    • Do not ramp HMS Q2 to zero, only go to 10A and then switch off power.
    • If any magnet trips, snapshot the associated power supply window and post it to hclog before resetting the interlock.
  • Monitor the beam position on target and call MCC immediately if you see large excursions (1 mm level).
  • Ramp rate lock instructions are here and added to checklist[5]

Bluejeans remote connection information

 To join the Meeting:
 https://bluejeans.com/861439920
 To join via Room System:
 Video Conferencing System: bjn.vc -or-199.48.152.152
 Meeting ID : 861439920
 To join via phone :
 1)  Dial:
   +1.408.740.7256 (US (San Jose))
   +1.888.240.2560 (US Toll Free))
   +1.408.317.9253 (US (Primary, San Jose))
   (see all numbers - http://bluejeans.com/numbers)
 2)  Enter Conference ID : 861439920

Useful Links

Target Polarization Angle Conventions

Be VERY clear here about the convention (beam-right/left/up/downstream) vs 0, 90, 180, 270)
for each run period.  More notes and references to log entries are better.

Polarization Calibration Data

ADD LIST OF CALIBRATION RUNS HERE
AND A TABLE OF RELEVANT CALIBRATION CONSTANTS TO USE FOR WHAT RUNS
Cell Calibration
Type
SHMS Angle
/degree
SHMS Momentum
/GeV
Field Direction
/degree
Calibration Results Status Oven Temperature
/°C
Log#
Tommy EPR-NMR 11 7.5 90 5.24 %/mV Done 200 3839939
Tommy EPR-NMR 18 5.6 270 5.04(1) %/mV Done 205 3843301
Tommy EPR-NMR 18 5.6 90 undone 205
Tommy EPR-NMR 11 7.5 270 undone 208
Tommy AFP Loss 14.5 6.4 90 1.25%/Sweep Done 200 3840618, 3840143, 3840618
Tommy EPR Amplifier
on/off NMR Signal
11 7.5 90 Done 200 3839962
Tommy Temperature Test 18 5.6 90 Done 205 3844089
Tommy AFP Loss 18 5.6 270 Done 205 3845255
Tommy Temperature Test 18 5.6 270 Done 205 3845336
Briana EPR-NMR 18 5.6 180 8.60 Done 205 3834036, 3834243
Briana EPR-NMR 18 5.6 270 5.6 Done 205 3835512, 3835587
Austin EPR-NMR 18 5.6 90 4.88 Done 200 3814717
Austin EPR-NMR 11 7.5 180 7.64 Done 210 3817200
Austin EPR-NMR 11 7.5 0 8.13 Done 210 3818761
Austin EPR-NMR 11 7.5 270 4.56 Done 210 3819882
Austin EPR-NMR 11 7.5 270 4.58 Done 210 3822296
Austin EPR-NMR 11 7.5 90 4.91 Done 210 3824907
Austin EPR-NMR 18 0 90 4.58 Done 210 3815895
Austin EPR-NMR 11 0 180 6.91 Done 210 3816724

Target Polarization Database

d2n Run Period

D2n Target Run List: https://docs.google.com/spreadsheets/d/16TrIUXVA-znVNUyEKy9FHaIlcvq-KXzu97k9nJowuE8/edit#gid=0

  • Junhao Chen is maintaining it

A1n Run Period

https://docs.google.com/spreadsheets/d/e/2PACX-1vSoTyDugD8wt3hKmBGZBxwUJTPHmUgyvGVJvM7N6qeVoNWhEXumCwwkPI7Sw9Asae3N4k9h5L0Y7qmw/pubhtml

  • Mingyu Chen will updating the runlist while Murchhana Roy will do the cross check.

Recent polarization measurements

PolHe3 Polarization Measurements






Misc Links